# Characterising and tackling thermally induced zero-drift in displacement measuring interferometry using temperature-controlled enclosure

### v1

@@ 4 Experimental investigations with the enclosure
We present several use cases where the characterization of (residual) thermal drifts is proven to provide useful qualitative results. The standard arrangement of the experiments, shown in Figure 3, was (with only a slight variation) used across most of the individual experiments.
The servo-controlled temperature in the enclosure allowed for two types of investigation. The temperature is kept stabilized for a prolonged period (days to weeks) so that the limits of the thermo-mechanical stability of individual components or more complex assemblies can be investigated. The complementary approach (that proved more useful) was to apply the thermal load in predefined cycles with linear ramp-up. With the bulk interferometers, the feasible temperature change rate was identified as K / hour. This value corresponds to quite a small power rating of the cooling relative to the enclosure volume. A faster heating/cooling rate led to more destabilized thermal conditions in the enclosure and, in turn, distorted measurement data.
The following use cases represent several experimental scenarios that address different research questions revolving around studying the thermal drifts in length measurement instrumentation. Combined together, they give an indicative overview of the range of testing and characterization possibilities.
@@ 4.1 Comparing variants of interferometric arrangement
In this experiment, we have compared two two-beam double pass interferometers (non-differential; a standard style used e.g. for coordinate positioning). One of the interferometers had the optical arrangement slightly modified so that the optical paths of the individual interferometer’s arms within the bulk optics of the interferometru were of equal length (so that one interferometer had the beams’ optical paths of different lengths, and the other had them of equal length). This modification was expected to reduce the thermal drifts and the objective of the test was to confirm that the modification of the beam paths will have desired effect.
The interferometers were fixed on standard kinematic mounts (Thorlabs) and aligned against silver mirrors. The interferometers were loaded with a temperature change of K with the cooling/heating rate of K/hour. The observed drifts, shown in Figure 4, were highly correlated to the induced thermal cycling () and improved more than twofold reduction. Nonetheless, the residual drift of the ”balanced” interferometer still came out significant, and it could be primarily attributed to thermal expansion of the opto-mechanics used since the interferometer are not differential.
@@ 4.2 Identifying faulty optical assembly
The aim of the test was to investigate drifts of a differential interferometer (modified arrangement[13]) produced with a modified assembly procedure. The interferometer was tested in the enclosure (without the active thermal control at that time) to investigate the drifts using the approach shown in Figure 3.
The interferometer did not show any signs of malfunctions (such as poor contrast, increased noise, presence of etalon effects, or quadrature signal distortion). The repeated tests in the enclosure indicated significant drifts (see the upper part in Figure 5) of hundreds of nanometres (over several weeks) instead of expected fluctuation at a nanometre level (see the lower part in Figure 5). After all other components were gradually replaced to rule out their contribution, the interferometer underwent partial disassembly. A loosened single corner of a retarder waveplate was revealed and identified as the source of drifts.
The conclusion is that zero-drift testing has a significant potential not only in characterizing the drifts but also in verifying the conformance of the interferometers (and generally any measuring instruments). The prior knowledge that is obviously necessary could come either from the previous experimental results or from analytical sources (simulations, calculations). The two significant benefits are that the technique could reveal faults otherwise difficult to detect and that the approach is simplistic enough to be useful for 100% verification of optical systems.
@@ 4.3 Improving mounting stability
Another measurement scenario demonstrates that the range of testing possibilities with the enclosure is not limited to the interferometer optics. When asked to characterize a pair of commercial interferometers, we used the standard experimental arrangement (Figure 3) to test both simultaneously. The interferometers were mounted in a vertical orientation according to the mechanical interface specified in the device documentation. Generic opto-mechanical components (Thorlabs) were used as the baseplates.
With the rough idea of the differential double-pass interferometer’s arrangement, we expected fairly pronounced thermal drift. The test results, however, revealed a significant global drift of several hundred nanometres (see the upper part in Figure 6). In several consecutive attempts, the fixture was gradually replaced with more robust components with enhanced triangular supports and additional side supports were added. With these measures applied, the observed drift was reduced by a factor of five with one of the interferometers and even more with the second one of the pair (see the lower part in Figure 6). Notably, the effect of the thermal load was still less significant than the residual (global) drift.
With the analysis of thermal influences presented before in mind, this type of test could provide a useful tool in the characterization and verification of complex (sub-) systems intended to become part of measurement instrumentation. It is also evident that the individual manifestations of thermal changes are principally superimposed and challenging to discern, and their mitigation requires careful consideration in multiple iterations.
@@ 4.4 Comparison of the assembly technology
The last experimental investigation we present is similar to the comparison of interferometers with the slightly varied arrangements (Section 4.1). This time, we loaded the enclosure with three interferometers of identical arrangement but differing in the assembly technique and, consequently, the glass material of the bulk components. The first interferometer has the components separately glued into the metal (aluminium) housing (cage assembly). In the second case, the optical components were cemented together (glue assembly), and the last interferometer was assembled using optical contacting (wringing). The first two interferometers were made of the N-BK7 and N-SF14 glasses (Schott), while the last one was made of fused silica. Instead of the standard optomechanics, the interferometers were fixtures on a titanium baseplate with additional support for the mirror’s kinematic mount. Due to the improved fixture, the global drift has been significantly reduced. The results of the test, displayed in Figure 7, indicate the temperature sensitivity is between nm.K and nm.K. The interferometer assembled by optical contacting exhibits the highest sensitivity because of the material (remember Table 1): fused silica has a significantly higher thermo-optic coefficient, so the influence on the optical lengths is the most pronounced. The sensitivity of the other two interferometers is comparable. What remains unaccounted for is the thermal expansion of the titanium baseplate, so that the inferred temperature sensitivity figures shouldn’t be considered absolute.
@@ 5 Final words
We have presented the achieved status in designing and developing a thermal-controlled enclosure dedicated to testing and characterising displacement measuring interferometers and their assemblies. The servo-loop temperature control achieves temperature stability of during a day and enables setting the temperature in the range of several kelvins. The enclosure could be used to keep the temperature stable over days and weeks and also makes possible to apply controlled thermal load to the devices under test.
@@
We demonstrated four use cases that used either of the capabilities for various characterization or testing tasks. The comparisons of zero-drift between similar interferometers (different assembly techniques, variant optical arrangement) provided valuable feedback for expanding the capabilities of and refining the interferometric measurement technology. Similarly, using the thermal load for intermittent testing provided feedback to redesign the interferometer’s mechanical fixtures, which significantly reduced the drifts. The enclosure’s functionality proved feasible for routine verification of the optical and optomechanical systems, where the testing under thermal load could help identify otherwise undetectable failures in the systems. Last but not least, the construction of the temperature-controlled enclosure and characterization of the thermal conditions inside could be transferred to the design of measurement instrumentation. The proper, intelligent and highly precise thermal management should be an intrinsic part of the design (e.g. [23]) not only in the top-grade instruments at national metrology laboratories – this an appealing approach to carrying out measurements under defined conditions, and simultaneously a necessary condition on the way to lower measurement uncertainty and an extended range of measurement scenarios.
@@
While the particular results we presented tend to be rather indicative than absolute, the message and contribution of this paper to our ultimate belief are demonstrative and methodological. The current state of the work opens both apparent questions and challenges (e.g. to fine-tune the temperature control servo loop, to develop bases for non-differential interferometers from low-expansion material) as well as the broader questions with overlap to precision engineering and measurement technology (e.g. investigation of the temperature distribution in the thermally controlled environment, because the temperature stability still relies on a single point measurement). Back to the interferometry, we also believe this work is crucial on the way towards further expansion of interferometric technology to routine operation in longer timeframes or in a less controlled environment – for instance, in the heart of a nuclear power plant[8].
@@ Acknowledgement
Technology Agency of the Czech Republic (FW03010687;TN02000020); European Commission (CZ.1.05/2.1.00/01.0017, LO1212, CZ.02.1.01/0.0/0.0/16_026/0008460); Czech Academy of Sciences (RVO:68081731)
@@
The authors declare no conflicts of interest.

### v2

@@ 4 Experimental investigations with the enclosure
In the following sections, we present particular experimental studies exploiting the (thermally-induced) zero-drift investigation to characterize or verify displacement-measuring interferometers.
The first experimental measurement investigates the (primarily temperature-induced) zero drift in two variants of single-ended interferometer arrangement and compares the result to the theoretically estimated figures. This experiment, described in section 4.1, demonstrates how to characterize the zero-drifts and thus, e.g. establish the upper boundaries for uncertainty assessment for interferometric measurements or allow for deterministic compensation of the effects. This is significant both in less-controlled environments and in longer time-frames – in both cases, both the temperature-induced and general zero-drift where the temperature effects are of more substantial importance.
Once the zero-drift behaviour is characterized for a particular interferometer, this knowledge is useful as a basis for (quality) testing or verification. This aspect is demonstrated in the second experimental study (section 4.2), where the experimental test revealed an assembly failure in the interferometer’s optics that would be otherwise hardly detectable.
Also, knowing the zero-drift of the interferometer’s optics, it is possible to extend the characterization to a more complex assembly. In section 4.3, we present how the thermal-load testing provided useful feedback in adjusting the mechanical fixture, i.e. it helped verify the system’s thermal stability.
Finally, thermal testing can help aid the design or manufacturing decision; in the case of interferometric optics, it is, for example, the assembly technology. In section 4.4, we present the investigation and comparison of three interferometers of identical geometry and dimension but with different assembly methods and compare the experimental results with theoretical predictions.
All the use cases represent several experimental scenarios that address different research questions revolving around studying the thermal drifts in length measurement instrumentation. Combined, they give an indicative overview of the range of testing and characterization possibilities.
@@ 4.1 Comparing variants of interferometric arrangement
In this experiment, we have compared two similar two-beam double pass single-ended (i.e. non-differential) planar interferometers, displayed in Figure 5 a) and b). This type of interferometer is commonly used, e.g. for coordinate positioning. The two variants have slightly different layouts of the optical paths of the two arms, and, as a result, the optical path difference within the interferometers’ optics is significantly different.
In the variant with two cube-corners (Figure 5 a)), the reference beam passes through the central beam-splitter twice, while the object beam does four passes and, additionally, it goes through the quarter-wave retarder (QWP) four times. Thus, the optical paths will differ in length by OPD , where is the nominal dimension of the beam-splitter and stands for the waveplate thickness. In contrast, in the other arrangement (Figure 5 b)), both arms pass the beam-splitter and either of the QWPs four times, so the beam paths within the optics are balanced.
Based on this consideration, we can estimate the optic thermal drift in the unbalanced interferometer (the drift presence of which has been mentioned decades ago [21]). The interferometers’s beam splitters were made of N-SF14 flint glass (Schott) and their nominal dimension was mm. The waveplates of crystal quartz had a thickness of mm. Thus, the beam paths differ by mm, so the optical path difference (incorporating thermal expansion and thermo-optic coefficient) would drift by nm per kelvin, so the total thermal sensitivity of the displacement measurement will be nm.K-1 (considering the double-pass layout). With the balanced arrangement, the thermal drifts are expected to reduce to zero.
We have recently carried out a similar study with two variants of a differential interferometer[14] and demonstrated significant improvement in the thermal resilience [9]. The study presented here, particularly the quantitative assessment of the thermal sensitivity, is significantly complicated because the interferometers are not of a differential arrangement. Consequently, the metrological loop is considerably longer and more complex, e.g. the physical distance between the interferometer is also susceptible to thermal expansion and adds up to the observed drifts.
In this experiment, the interferometers were mounted on dedicated holders made of low expansion Titanium (depicted in Figure 5 c)) and aligned against planar mirrors (Zerodur + dielectric coating). While it is a straightforward choice to use the low-expansion materials for the optomechanics, there is also a significant drawback: the same materials also exhibit low thermal conductivity, so in turn, it poses a further limit on the temperature loading bandwidth.
To estimate the contribution of the mounting, we assume both the mirror and the interferometer to be fixed to the holder (bed) at the centre of gravity (marked with the dash-dotted lines). Then we accounted for the relevant material amount of the mirror (denoted , mm), the bed (mm) and the interferometer (mm). The total contribution to the temperature sensitivity was approximately nm.K-1. Nonetheless, the contributions of individual terms reaches almost a hundred nanometers, so we expect quite high uncertainty associated with this estimation.
For this experiment, we used the temperature change of K with the cooling/heating rate of K/hour. The experimental results are shown in Figure 6. The interferometers’ displacement reading have a residual linear drift subtracted: nm.hr-1 for the unbalanced interferometer and nm.hr-1 for the balanced one. The observed drifts were highly correlated to the induced thermal cycling (), but the observed hysteresis indicates the presence of either a residual drift component or too steep temperature change.
Subtracting the contribution of the holder, the total observed thermal sensitivity is nm.K-1 for the unbalanced interferometer and nm.hr-1 for the balanced one. These figures correspond well to the estimations (coindcidence nm.K-1). Nonetheless, we expect the associated uncertainty to be considerable (probably nm.K or more) for a number of reasons, where the imperfection of the mounting thermal response model (with several questionable parameter values, such as the effective mounting pivot points) would be the dominant one. Also, the glue layers between the optical elements of the interferometer and the dielectric coatings remain unaccounted for. On the other hand, with the small beam path on the air (mm), the influence of the refractive index of air would be less than nm.K-1 combined.
@@ 4.2 Identifying faulty optical assembly
The aim of the test was to verify that a manufactured sample of a double-pass planar differential interferometer [9]. We tested whether it has the expected thermal sensitivity compared to the previously observed performance we achieved with the reference interferometer of the same construction (see the top plot in Figure 7). With this interferometer, we observed the drifts in the range smaller than nm peak-to-peak over hour timeframe with the temperature fluctuation of K (the measurement did not take place in the enclosure). The residual fluctuations were attributed to the polarization fluctuations in the fibre feed.
The interferometer under test was put in the enclosure (without the active thermal control at that time) to investigate the drifts using the standardized experimental testbed and protocol (Figure 3). The interferometer was mounted flat onto a raised platform with a silver mirror in kinematic mount, placed mm apart from the interferometer’s front side.
The tested interferometer did not show any signs of malfunctions (such as poor contrast, increased noise, presence of etalon effects, or quadrature signal distortion). The test in the enclosure revealed significant drifts (see the lower part in Figure 7) of hundreds of nanometres (over several weeks; achieved temperature stability was K) instead of expected fluctuation at a nanometre level. After all other components were gradually replaced in the measurement chain (laser, fibre feed, receiver, mirror with holder) to rule out their contribution, the interferometer underwent partial disassembly. A loosened single corner of a retarder waveplate was revealed and identified as the source of drifts.
The moral of this experiment is that zero-drift testing has a significant potential not only in characterizing the drifts but also in verifying the conformance of the interferometers (and generally any measuring instruments). The prior knowledge that is obviously necessary could come either from the previous experimental results or from analytical sources (simulations, calculations). The two significant benefits are that the technique could reveal faults otherwise difficult to detect and that the approach is simplistic enough e.g. to be useful for 100% verification of optical systems during manufacturing.
@@ 4.3 Improving mounting stability
Another measurement scenario demonstrates that the range of testing possibilities with the enclosure reaches beyond the interferometer optics itself. When asked to characterize a pair of commercial interferometers, we used the standard experimental arrangement to test both interferometers simultaneously. The interferometers were mounted vertically according to the mechanical interface specified in the device documentation. Generic opto-mechanical components were used as the baseplates.
With the rough idea (the manufacturer did not disclose the detailed arrangement) of the differential double-pass interferometer’s arrangement, we expected fairly pronounced thermal drift. The test results, however, revealed a significant global drift of several hundred nanometres (see the upper part in Figure 8). In several consecutive attempts, the fixture was gradually replaced with more robust components with enhanced triangular supports and additional side supports were added. With these measures applied, the observed drift was reduced by a factor of five with one of the interferometers and even more with the second one of the pair (see the lower part in Figure 8). Notably, the effect of the thermal load was still less significant (which was a valuable outcome of the test) than the residual (global) drift.
With the analysis of thermal influences presented before in mind, this type of test could provide a valuable tool in the characterization and verification of complex (sub-) systems intended to become part of measurement instrumentation. It is also evident that the individual manifestations of thermal changes are principally superimposed and challenging to discern, and their mitigation and fine-tuning requires careful consideration in multiple iterations. The zero-drift analysis can aid the effort and provide necessary feedback during the process.
@@ 4.4 Comparison of the assembly technology
The last experimental investigation we present is similar to comparing interferometers with the slightly varied arrangements (Section 4.1). This time, we loaded the enclosure with three interferometers of identical arrangement (the unbalanced planar interferometer, depicted in Figure 5 a)) but differing in the assembly technique and, consequently, the glass material of the bulk components. The goal was to compare the thermal sensitivity to otherwise identical interferometers experimentally.
The first interferometer has the components separately glued into the metal (aluminium) housing (cage assembly). In the second case, the optical components were cemented together (glue assembly), and the last interferometer was assembled using optical contacting (wringing). The first two interferometers were made of the N-BK7 and N-SF14 glasses (Schott), while the last one was made of fused silica. As mentioned before, the estimated thermal sensitivity of the N-BK7+N-SF14 interferometers were estimated to be nm.K-1. Fused silica has a lower coefficient of thermal expansion, nonetheless, it has a significantly higher thermo-optic coefficient, so the temperature influence on the optical pathlengths more pronounced, and the total thermal sensitivity is nm.K-1.
The interferometers were mounted on a titanium holder (Figure 5 c)) with a silver mirrors (instead of the rectangular Zerodur mirrors used in Section 4.1). This change and using a larger separation between the mirror and the front face of the interferometer caused the contribution to the zero drift to rise to approximately nm.K-1.
The results of the test, displayed in Figure 9, indicate the total temperature sensitivity of the assembly (interferometer + mirror + holder) is between nm.K and nm.K. Similarly to the experiment from Section 4.1, the observed drifts were well correlated to the induced thermal cycling (), and noticeable hysteresis of (a symmetric nature) indicating too steep temperature change was observed.
Taking the contribution of the holder, we get approximate coefficients as follows: cage assembly nm.K, glue assembly nm.K, optical contacting nm.K. All figures coincide with the theoretical estimation within nm.K. With the cage assembly, the physical separation of the components (unlike the other two assemblies) probably makes the interferometer partially self-compensated against the thermal drifts.
@@ 5 Final words
We have presented the achieved status in designing and developing a thermal-controlled enclosure dedicated to testing and characterizing displacement-measuring interferometers and their assemblies. The servo-loop temperature control achieves temperature stability of mK during a day and enables setting the temperature in the range of several kelvins. The enclosure could be used to keep the temperature stable over days and weeks, and it could also be used to apply a controlled thermal load to the devices under test.
@@
We demonstrated four use cases that used either of the capabilities for various characterization or testing tasks. The comparisons of zero-drift between similar interferometers (different assembly techniques, variant optical arrangement) provided valuable feedback for expanding the capabilities of and refining the interferometric measurement technology. Generally, this type of study can aid the development process, provide experimental verification of the instrumentation design, and provide critical input to the uncertainty assessment. Similarly, using the thermal loading for intermittent testing provided feedback for redesigning the interferometer’s mechanical fixtures, significantly reducing the drifts and demonstrating the feasibility of using the zero drift investigations for iterative improvement and verification in the development of metrological instrumentation. Similarly, the methodology proved feasible for routine verification of the optical and optomechanical systems, where the testing under thermal load could help identify otherwise undetectable failures in the systems.
The results from the presented investigations also have a potential impact on particular applications. The observed magnitude of thermally-induced zero drifts might interfere with the accuracy requirements in the long-range coordinate positioning (mm), either in air (nanometrology) or in vacuum (precise microscopy, e-beam lithography), also in 1-D length calibration instrumentation for secondary- or shop-floor metrology in the range up to m (the influence of the refractive index would prevail above this level).
Last but not least, the construction of the temperature-controlled enclosure and characterization of the thermal conditions inside it could be transferred to the design of measurement instrumentation: Within the future development of dimensional metrology instrumentation, proper, intelligent and highly precise thermal management should become an intrinsic part of the design (e.g. [27]) not only in the top-grade instruments at national metrology laboratories – this an appealing approach to carrying out measurements under defined conditions, and simultaneously a necessary condition on the way to lower measurement uncertainty and an extended range of measurement scenarios.
@@
While the particular quantitative results we presented might be considered indicative rather than absolute, the message and contribution of this paper to our ultimate belief are demonstrative and methodological. The current state of the work opens both apparent questions and challenges as well as the broader questions with overlap to precision engineering and measurement technology (e.g. investigation of the temperature distribution in the thermally controlled environment, because the temperature stability still relies on a single point measurement). Back to the interferometry, we also believe this work is crucial on the way towards further expansion of interferometric technology to routine operation in longer timeframes or in a less controlled environment – for instance, in the heart of a nuclear power plant[8].
@@ Acknowledgement
Technology Agency of the Czech Republic (FW03010687;TN02000020); European Commission (CZ.1.05/2.1.00/01.0017, LO1212, CZ.02.1.01/0.0/0.0/16_026/0008460, CZ.02.01.01/00/22_008/0004649); Czech Academy of Sciences (RVO:68081731)
@@
The authors declare no conflicts of interest. While preparing this work, the authors used the Grammarly tool (www.grammarly.com) to correct the language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.